A slab consists of two parallel layers of two different materials of same thickness having thermal conductivities k1 and k2. The equivalent conductivity of the combination is

Grade:11

2 Answers

Bhagyesh
7 years ago
H = H1 + H2 for slabs in parallel. ∆T.KA/L = ∆T.K1.A/2L + ∆T.K2.A/2LHere for k1 and k2 we take A/2. That`s imp. So simplify and get K = K1 + K2 / 2
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

H = H1 + H2 for slabs in parallel.
∆T.KA/L = ∆T.K1.A/2L + ∆T.K2.A/2L
Here for k1 and k2 we take A/2.
So simplify and get K = (K1 + K2) / 2

Thanks and Regards